# Event-based Photometric Bundle Adjustment

### v1

@@ 1 Introduction
Figure 1: EPBA jointly refines the camera rotations and panoramic intensity map by minimizing a photometric error. Each event contributes an error term by computing the difference of the intensities at two map points.
@@ 3.1 Event Generation Model (EGM)
where the event polarity $s_{k}\in\left\{+1,-1\right\}$ indicates the sign of the change [13], and $\Delta t_{k}$ is the time elapsed since the last event at the same pixel $\mathbf{x}_{k}$. This is illustrated in Fig. 3(a): induced by the camera motion, the brightness at pixel $\mathbf{x}_{k}$ decreases from white (snow) at $t_{k}-\Delta t_{k}$ to gray (sky) at time $t_{k}$, thus producing a negative event $e_{k}$.
EGM in the scene. Assuming a 3D log-intensity map $M$ of the scene, that light travels without attenuation between the scene and the image plane, and that the event is due to the relative motion between the camera and the scene, one may cast the equation on the image plane (1) into an equation in terms of map intensities:
$\Delta L\equiv M(\mathbf{S}(t_{k}))-M(\mathbf{S}(t_{k}-\Delta t_{k}))=s_{k}C,$ (2)
where map points $\mathbf{S}(t_{k}-\Delta t_{k}),\mathbf{S}(t_{k})$ project on the same camera pixel, $\mathbf{x}_{k}$, at times $t_{k}-\Delta t_{k}$ and $t_{k}$, respectively. Hence, the EGM (2) naturally associates each event with two map points. This is illustrated in Fig. 3(b): As the camera moves, the pixel $\mathbf{x}_{k}$ scans different parts of the scene. Letting $M(\mathbf{S})$ represent the intensity of the scene map at surface point $\mathbf{S}$, the intensity decreases from the white snow value $M(\mathbf{S}(t_{k}-\Delta t_{k}))$ to the gray sky intensity $M(\mathbf{S}(t_{k}))$ (here we assume there is a sphere of infinite radius that provides values for non-surface objects like the sky). The change of intensity between two map points describes the negative event $e_{k}$. This is also illustrated in Fig. 1: as the camera rotates, the ray through pixel $\mathbf{x}_{k}$ (of event $e_{k}$) touches two map points (with panoramic coordinates $\mathbf{p}(t_{k}-\Delta t_{k})$ and $\mathbf{p}(t_{k})$), whose intensity difference (akin (2)) describes $e_{k}$.
@@ 3.3 Formulation for a Rotational Event Camera
The event pixel location $\mathbf{x}$ is back-projected to $\mathbf{X}\in\mathbb{R}^{3}$, rotated $\mathbf{X}^{\prime}=\mathtt{R}(t)\mathbf{X}\in\mathbb{R}^{3}$, and projected onto a (panoramic) map point $\mathbf{p}$. Though it is possible to define a map on the sphere, $\mathbf{X}^{\prime}\equiv\mathbf{S}$, for simplicity we consider a 2D rectangular representation (e.g., equirectangular projection $\pi$). Compounding transformations, $\mathbf{x}$ is transferred (i.e., warped) to the map point $\mathbf{p}$ according to the camera orientation $\mathtt{R}(t)$, the intrinsic calibration $\mathtt{K}$ and the type of projection model $\pi$ used to represent the map: $\mathbf{x}\mapsto\mathbf{p}$, i.e.,

### v2

# {textblock}

@@ 1 Introduction
Figure 1: EPBA jointly refines the camera rotations and panoramic intensity map by minimizing a photometric error. Each event contributes an error term by computing the difference of the intensities at two nearby map points.
@@ 3.1 Event Generation Model (EGM)
Let $L$ represent the logarithmic intensity of light at the sensor [17], and $M$ be the logarithmic intensity (radiance) on the surface $\mathbf{S}\subset\mathbb{R}^{3}$ of objects in the scene. We formulate the well-known idealized version of the event-triggering condition and then rewrite it in terms of the scene elements.
@@
where the event polarity $s_{k}\in\left\{+1,-1\right\}$ indicates the sign of the change [13], and $\Delta t_{k}$ is the time elapsed since the last event at the same pixel $\mathbf{x}_{k}$. This is illustrated in Fig. 3(a): induced by camera motion, the brightness at pixel $\mathbf{x}_{k}$ decreases from white (snow) at $t_{k}-\Delta t_{k}$ to gray (sky) at time $t_{k}$, thus producing a negative event $e_{k}$.
EGM in the scene. Assuming ($i$) that the event is due to the relative motion between the camera and the scene and ($ii$) that there is a geometric and photometric model of the scene, we may cast (1) from the image plane into an equation in terms of the scene elements. Let $\mathbf{S}(\mathbf{x},t)\in\mathbb{R}^{3}$ be the surface point that projects onto pixel $\mathbf{x}$ at camera viewpoint $\mathtt{P}(t)$ (classically, $\mathbf{S}(\mathbf{x},t)$ is the point in space where the ray back-projected through pixel $\mathbf{x}$ from camera viewpoint $\mathtt{P}(t)$ intersects the surface of objects in the scene). As photometric model, assume that light travels without attenuation between a scene point $\mathbf{S}$ and its projection $\mathbf{x}$ on the image plane [56], that is, the intensity of the scene map is $M(\mathbf{S}(\mathbf{x},t))\doteq L(\mathbf{x},t)$. Then, $\mathbf{S}(\mathbf{x}_{k},t_{k})$ and $\mathbf{S}(\mathbf{x}_{k},t_{k}-\Delta t_{k})$ are the scene points that project on the same camera pixel, $\mathbf{x}_{k}$, at times $t_{k}-\Delta t_{k}$ and $t_{k}$, respectively, and the intensities in (1) transfer to these 3D points, yielding an equation in terms of map intensities:
$\Delta L\equiv M\bigl{(}\mathbf{S}(\mathbf{x}_{k},t_{k})\bigr{)}-M\bigl{(}\mathbf{S}(\mathbf{x}_{k},t_{k}-\Delta t_{k})\bigr{)}=s_{k}C.$ (2)
Hence, (2) naturally associates each event with two map points (whose intensity difference is $s_{k}C$). This is illustrated in Fig. 3(b): as the camera moves, the pixel $\mathbf{x}_{k}$ scans the scene, with the radiance decreasing from the white snow value $M(\mathbf{S}(\mathbf{x}_{k},t_{k}-\Delta t_{k}))$ to the gray sky $M(\mathbf{x}_{k},\mathbf{S}(t_{k}))$ (assume there is a sphere of infinite radius that provides values for non-surface objects like the sky [56, 57]). The path a pixel takes across the scene is effectively described by the concatenation of segments. These segments are defined between pairs of consecutive points: the start and end points of each segment correspond to the time of the previous event $t_{k}-\Delta t_{k}$ and the time of the current event $t_{k}$ (at the same pixel $\mathbf{x}_{k}$), respectively.
It is also illustrated in Fig. 1: as the camera rotates, the ray through pixel $\mathbf{x}_{k}$ (of event $e_{k}$) touches two map points (with panoramic coordinates $\mathbf{p}(t_{k}-\Delta t_{k})$ and $\mathbf{p}(t_{k})$), whose intensity difference (akin (2)) describes $e_{k}$.
@@ 3.3 Formulation for a Rotational Event Camera
The event pixel location $\mathbf{x}$ is back-projected to $\mathbf{X}\in\mathbb{R}^{3}$, rotated $\mathbf{X}^{\prime}=\mathtt{R}(t)\mathbf{X}\in\mathbb{R}^{3}$, and projected onto a (panoramic) map point $\mathbf{p}$. Though it is possible to define a map on the sphere, $\mathbf{X}^{\prime}(t)\equiv\mathbf{S}(\mathbf{x},t)$, for simplicity we consider a 2D rectangular representation (e.g., equirectangular projection $\pi$). Compounding transformations, $\mathbf{x}$ is transferred (i.e., warped) to the map point $\mathbf{p}$ according to the camera orientation $\mathtt{R}(t)$, the intrinsic calibration $\mathtt{K}$ and the type of projection model $\pi$ used to represent the map: $\mathbf{x}\mapsto\mathbf{p}$, i.e.,